### v1

@@ Abstract
For a given finite (Galois) field $F\equiv\mathbb{F}_{q}$, of size $|F|=q=p^{m}$ and prime characteristic $p$, an $F$-linear code of length $n$ and dimension $k$ is a linear space ${\cal C}\subseteq F^{n}$ of dimension $k$. Any set of linearly-independent rows of a generator matrix $G$ forms a basis of the code ${\cal C}\equiv{\cal C}_{G}$ generated by $G$; the code dimension is $k=\mathop{\rm rank}G$. The dual code ${\cal C}^{\perp}\equiv{\cal C}_{G}^{\perp}$ of dimension $n-k$ is formed by all vectors in $F^{n}$ orthogonal to non-zero vectors in ${\cal C}$ or, equivalently, to the rows of $G$, a check matrix of the code ${\cal C}_{G}^{\perp}$.
@@
The usual inner product in $F^{\ell}$ is related to the group trace with the help of a linear map[37] ^$\,\boldsymbol{\widehat{\ }}:F[G]\to F[G]$,
@@
${\boldsymbol{a}}\cdot{\boldsymbol{b}}\equiv\sum_{g\in G}a_{g}b_{g}=\mathop{\rm Tr}\nolimits_{G}(\widehat{a}b)=\mathop{\rm Tr}\nolimits_{G}(b\widehat{a}).$ (4)
@@
′′′$\partial^{\prime\prime\prime}(a\otimes b)\equiv\partial^{\prime}a\otimes b+(-1)^{i}a\otimes\partial^{\prime\prime}b,$ (11)
@@
The homology groups of the product $\mathcal{C}={\cal A}\times{\cal B}$ are isomorphic to a simple expansion in terms of those of ${\cal A}$ and ${\cal B}$ which is given by the Künneth formula,
@@ III.1 Multi-block complexes and associated quantum codes
With Eq. (20), it is easy to verify that the resulting matrices contain only three blocks, $A\equiv A_{1}\otimes I_{B}\otimes C_{1}$, $B\equiv I_{A}\otimes B_{1}\otimes I_{C}$, and $C\equiv I_{A}\otimes I_{B}\otimes C_{1}$, and can be explicitly written as
@@
[…]
@@
For a $D$-dimensional MBC complex with boundary matrices $Q_{j}$, $1\leq j\leq D$, it is easy to verify that all $D$ blocks are used in an equivalent fashion, i.e., all chain complexes constructed from the same $D$ commuting blocks taken in different orders are equivalent (can be obtained from each other by block row and column permutations). In particular, matrix $Q_{j}$ has $j$ non-zero blocks in each column and $D+1-j$ non-zero blocks in each row. Furthermore, in each pair of matrices $Q_{j}$ and $Q_{j+1}^{T}$ with $m_{j}={D\choose j}$ column blocks, $1\leq j<D$, each pair of matching column-blocks contains a full set of blocks $\{A_{i}\}$, exactly $j$ such blocks in $Q_{j}$ and the remaining $D-j$ (transposed) blocks in $Q_{j+1}^{T}$.
@@
$[\mathop{\mathrm{M}}\nolimits(a)]_{\alpha,\beta}\equiv\sum_{g\in G}a_{g}\delta_{\alpha,g\beta},$ (58)
where group elements $\alpha,\beta\in G$ are used to index rows and columns, and $\delta_{\alpha,\beta}=1$ if $\alpha=\beta$ and 00 otherwise is the Kronecker delta. It is easy to verify that for group elements $g\in G$, matrices $\mathop{\mathrm{M}}\nolimits(g)$ form the regular $F$-linear representation of $G$.
@@ III.3 Equivalent AMC codes
Therefore, we can always “rescale” the elements in the original set, and choose $a_{j}=1+\ldots$, $j\leq D$. Further, we can try to construct a $D$-local complex in $\mathbb{Z}^{D}$, e.g., by choosing group generators in the support of each element, e.g., $a_{j}=1+x_{j}+(\text{other monomials})$. Clearly, the second transformation changes the group representation and the basis vectors associated with the lattice.
For example, with weight-two elements $a_{j}=1+x_{j}$, $1\leq j\leq D$, we get a complex locally equivalent to $D$-dimensional QHP codes. Assuming the entire group can be generated this way [^3] Otherwise, the spaces in an AMC complex can be decomposed onto a direct sum of subspaces corresponding to cosets of the subgroup $G_{X}=\langle\{x_{j}\}\rangle$ in the original group $G$; see Sec. IV-C in Ref. 28. In the abelian case these complexes are permutation-equivalent to a complex in the subgroup $G_{X}\leq G$.,
@@
we get a $D$-complex locally equivalent to $D$-dimensional toric codes. With a level-$j$ quantum CSS code, for errors with irreducible weight smaller than the injectivity radius $R$, this guarantees identical syndrome weights, that is, identical confinement profiles[22] up to $R$. We also expect the code distance to grow as $d={\cal O}(R^{\alpha})$, with the exponent $\alpha=\min(j,D-j)$ expected for level-$j$ $D$-dimensional toric codes, see Statement 4 below. Further, we can use a measuring circuit designed for a regular level-$j$ $D$-dimensional toric code: its validity is guaranteed as a local property. [^4] Note, however, that the circuit distance need not always coincide with that of the original code. As an example, while any single-ancilla measurement circuit would work for a toric code[59], rotated surface codes require a carefully designed $\sf Z$-$\sf N$ addressing scheme for fault-tolerance[51].
@@
As an immediate consequence, the corresponding homology groups have orders $\left|H_{j}({\cal Q})\right|={D\choose j}\kappa$. Previously, the same or equivalent results were derived for the special case of $D=2$ AMC codes (quasi-cyclic GB codes), see Refs. [24, 44]. Also note a similarity with the corresponding equations for $D$-dimensional HGP codes. The main difference with Eqs. (18) and (19) is the linear scaling with the group order $\ell$, which is why the AMC construction gives much smaller codes.
@@
Let $G$ be a finite abelian group presented with the list of generators $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{m})$, and $K\equiv G^{\times D}$ its $D$-fold tensor product with itself, with the corresponding list of generators $\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{D}\}$. Given a checkable[37] abelian group-algebra code in $F[G]$ of length $n$, dimension $k$, and distance $d$, with the orthogonal element $a\equiv a(\boldsymbol{x})$, consider the group algebra elements $a_{i}\equiv a(\boldsymbol{x}_{i})\in F[K]$, $1\leq i\leq D$, and the corresponding $D$-complex ${\cal Q}=\mathop{\text{\small\sc AMC}}\nolimits(a_{1},a_{2},\ldots,a_{D})$. Its $j$-th subspace has dimension $\dim({\cal Q}_{j})={D\choose j}n^{D}$, while the corresponding homology and co-homology groups have the same dimension $k_{j}={D\choose j}k^{D}$ and homological distances $d_{j}=d^{j}$ and $\tilde{d}_{j}=d^{D-j}$, respectively.
In the special case of symmetric even-$D$ AMC codes at level $j=D/2$, this gives the same rate as the original code, and the distance $d^{D/2}$. Thus, a finite rate and a linear distance scaling in the original code family would give a finite rate and the distance scaling as a square root of the code length, similar to GB codes. While we are not aware of a proof that good linear abelian group-algebra codes exist (known results include good code families based on non-abelian dihedral groups $D_{2m}$, random quasi-abelian codes [46], and non-linear cyclic codes[47] that attain the Gilbert-Varshamov bound), already the classical results on the cyclic codes[48, 49] are sufficient to show the existence of AMC codes with finite relative distances and power-law distance scaling, with the exponent arbitrarily close to $1/2$.
@@
Another parameter relevant in a fault-tolerant setting is the syndrome distance $d_{\rm S}$, the minimum weight of a non-zero syndrome vector. An obvious upper bound is the minimum column weight in the check matrix. A boundary operator $Q_{j}$ in an AMC $D$-complex ${\cal Q}$ has $j$ or $D+1-j$ non-zero blocks in each block column or row, respectively. If we use group algebra elements of the same weight $w$, this gives uniform column weights $jw$ and row weights $(D+1-j)w$. For a code $\mathop{\rm CSS}\nolimits(Q_{j},Q_{j+1}^{T})$ in the space ${\cal Q}_{j}$ of such a $D$-complex, this gives a simple upper bound on the syndrome distance,
@@ IV.1 Family of $4$-dimensional rotated toric codes
Specifically, we constructed such codes for $7\leq\ell\leq 30$ and corresponding block lengths $n=6\ell$ (four distinct polynomials cannot be found for $72\leq\ell<$ 7). Generating polynomials and parameters of smallest codes for each minimum distance $d$ are listed in Tab. 1. All constructed codes have the same dimension $k=6$ and the same characteristic polynomial $h(x)=1+x$. For comparison, in Tab. 1 we also list parameters of two conventional $4D$ toric codes, which require larger block lengths to achieve the same distances.
In addition to conventional code parameters and the defining polynomials, Tab. 1 shows the confinement profile[22], a sequence of minimum syndrome weights corresponding to errors of increasing irreducible weights $w\geq 1$. These are important for single-shot error correction in the presence of syndrome measurement errors. Indeed, for a conventional level-2 4D toric code, the unobservable errors are closed surfaces formed by plaquettes, while such a surface with a boundary is observable, where edges along the surface boundary correspond to flipped syndrome bits. With the lattice of size $L^{4}$, the perimeter is expected to scale as $\mathcal{O}(w^{1/2})$, for error weights $w\leq w_{\rm max}=\mathcal{O}(L^{2})$, with the first five terms in the large-$L$ sequence $4,6,6,8,8$. The confinement profiles of the codes in Tab. 1 (computed using the dist_m4ri program [50]) follow this sequence, as long as the corresponding minimum-perimeter configuration does not contain a pair of edges identified by a lattice vector. This is substantially different from $3+3$ GB codes previously studied in Ref. 25, where for all designed codes the confinement profiles (starting with $3,4,3$) are increasing only for irreducible error weights $w\leq 2$.
@@
Table 1: Smallest $D=4$ AMC2 qubit codes (rotated $4D$ toric codes) with increasing distances $d$, constructed from cyclic groups $C_{\ell}$ with $\ell\leq 30$ (above the line) and two conventional $4D$ toric codes (below the line). Confinement profile [22] (“confin”) column lists the minimum syndrome weights for irreducible errors of weights from $1$ to $5$, except for codes with $\ell=7$ and $\ell=2^{4}$ where the biggest irreducible errors have weights $4$ and $3$, respectively.
@@ IV.2 Circuit design
As an example, after removing the first row block from the CSS matrices $H_{X}=R_{2}$ and $H_{Z}=R_{3}^{T}$ in Eqs. (45) and (52), we get (we also dropped the signs assuming the binary field)
@@
where the subscripts next to block names label the columns and the circumflex “hat” accent indicates transposed blocks. Focusing on bigger $3\times 3$ blocks, we have a two-block structure similar to that in lifted-product or GB/2BGA codes, with two and four (”$2+4$”) non-zero entries in each subrow. The circuits designed[25] for $2+4$ 2BGA/GB codes, a generalization of the $\sf Z$-$\sf N$ addressing scheme[51] for rotated surface codes, also work in the present case. Namely, a valid time-optimal circuit is obtained if we address the qubits corresponding to the two terms in the $D$ blocks at steps 1 and 6 (first and last), with the opposite order for $H_{X}$ and $H_{Z}$, while addressing the qubits from the remaining blocks in the intervening time steps. Specifically, we used the following addressing scheme:
@@
Here each column corresponds to a circuit time step, and each row to a block row of the matrices (67), (71), with secondary superscripts labeling the location corresponding to a free term (“0”) and the other monomial (“1”). Circuits corresponding to the original matrices with 2nd, 3rd, or 4th row removed were constructed identically, with the help of column permutations rendering the matrices in the form (67), (71), up to block labels.
Based on this addressing scheme for a single measurement round, we designed three types of four-round syndrome measurement circuits, $1$, 1212, and 1234555To guarantee rank preservation after removal of the 1st block row (which is important for the $1$ cycle), guided by Lemma 6, the polynomials were ordered so that $a_{4}=1+x$ [which gives block $D$ in Eqs. (45) and (52)].. Here a digit stands for the row omitted in each particular round of measurement, and we constructed detector events matching pairs of closest available measurement outcomes, either from neighboring or next-nearest-neighboring rounds. Overall, all our circuits have fixed duration of 9 measurement rounds, including two four-round cycles and a final measurement of all data qubits in $Z$ or $X$ basis, depending on the initial state of the data qubits, |$\ket{0}^{\otimes n}$ or |$\ket{+}^{\otimes n}$. The results of the final measurement are combined to construct the final set of detector events (thus computed syndrome bits are exact), and also the logical observables which allow to verify whether the decoding was successful.
@@
For accurate decoding, one needs to know probabilities of individual faults in the circuit and their effects on the detectors and the logical observables. This information for a given circuit is computed automatically by Stim, which also aggregates the faults producing the identical effect, and can be exported as a detector error model (DEM) file[53]. Information about aggregated faults in a DEM file can be regarded as columns of stabilizer-generator $H_{X}$ and logical $L_{X}$ matrices (with probabilities for each column also known) for an asymmetric quantum CSS code. In particular, the distance $d_{Z}$ of this code is identified as the circuit distance. We used vecdec program [54] to export the matrices from the DEM files associated with individual circuits, and dist_m4ri program to compute the corresponding distances (with the distances $d\leq 6$ computed exactly with a greedy enumeration algorithm, and upper bounds for larger distances using a random information set algorithm[55], exact with a high probability).
@@
Circuit simulations have been done with Pauli-frame circuit simulator Stim [53] written by Craig Gidney, and the decoding using vecdec program [54]. Specifically, before attempting belief-propagation (BP) decoding, the program tries to match the syndrome data against a pre-computed list of small-weight error clusters; such a simple pre-decoder gives a substantial acceleration, especially at smaller error probabilities. If pre-decoding fails, the corresponding clusters are discarded, and a BP decoder with serial schedule, using both average and instantaneous logarithmic likelihood ratios (whichever converges first) is engaged, with up to 10 steps. Finally, if that fails, the logarithmic likelihood ratios (LLRs) are transmitted to the OSD-1 decoder. A sample of circuit simulation data for codes $[[42,6,4]]$, [66,6,6]][66,6,6]][ 66 , 6 , 6 ] ], and $[[96,6,8]]$ (see Tab. 1), with the standard error model described earlier and the “1212” syndrome-addressing cycle, is shown on a threshold-style plot in Fig. 1. To speed up the decoding, we simplified the circuit by dropping all “minority” detector events [^6] Unlike with the BP+OSD package[60, 61] whose performance may be degraded by additional detector events, see in Appendix A of Ref. 62, for the decoder we used, removing minority detector events increases logical error rates, e.g., from $0.20\%$ to $0.36\%$ at $p=0.1\%$ with the $[[42,6,4]]$ code and the “1212” circuits. Asymptotic slope at small $p$ consistent with $d/2$ (where $d$ is the code distance) is in agreement with the circuit distances computed independently. The (pseudo)threshold position (crossing point) is at $p_{c}\approx 1.1\%$, in approximate agreement with the circuit simulation results for similar codes in Ref. 34. We note that our value is (marginally) higher than the values for toric and surface codes under the same error models.
@@
Figure 1: Logical error rate as a function of the circuit noise parameter $p$. Results for short even-distance codes $[[42,6,4]]$, [66,6,6]][66,6,6]][ 66 , 6 , 6 ] ], and $[[96,6,8]]$ (see Tab. 1) are shown. Inset shows the success probability $1-p_{L}$ as a function of $p$ in the vicinity of the crossing point(s) at $p\approx 1.1\%$
@@
In addition, to verify the “single-shot” properties of the constructed codes, we have also tried sequential one-step sliding window decoding[18, 57, 23], with windows of different sizes. Details of our implementation are given in Ref. 25.
Some simulation results for the code $[[42,6,4]]$ are shown in Fig. 2, which shows the logical error rates $p_{L}$ as a function of the decoding window sizes. These results correspond to data points shown with open symbols, with three sets corresponding to different addressing cycles as indicated in the caption. While all cycles give nearly identical error rates with full-window ($T=9$) decoding, the “$1$” cycle (which discards most of the redundant stabilizer generators) shows the highest error rates with single-shot ($T=1$) and two-shot ($T=2$) decoding protocols. In comparison, the circuits designed with “1234” and “1212” cycles give nearly the same decoding accuracy with $T=2$ (two-shot) and $T=9$ (full-window) decoding. Note that only during the rounds labeled with “1” a full-rank set of stabilizer generators is addressed; this explains a slightly better accuracy of the “1212” measurement cycle compared to “1234”.
Figure 2: Logical error rate for the $[[42,6,4]]$ AMC code using full detector error model (both $X$ and $Z$ detector events) as a function of the decoding window size $T$; $T=9$ corresponds to full-block decoding, while the window sizes $T<9$ correspond to a sliding window decoding protocol, see text. The curves labeled in the caption as “$1$”, “1234”, and “1212” correspond to different measurement cycles as explained in the text. Single-shot decoding ($T=1$) is the least accurate as a single fault can sometimes lead to a decoding error. Two-shot decoding, $T=2$, with the “1234” and “1212” measurement cycles is nearly as accurate as full-block decoding, $T=9$. This is unlike the “$1$” measurement cycle where most of the redundant checks are discarded.

### v2

@@ Abstract
For a given finite (Galois) field $F\equiv\mathbb{F}_{q}$, of size $|F|=q=p^{m}$ and prime characteristic $p$, an $F$-linear code of length $n$ and dimension $k$ is a linear space ${\cal C}\subseteq F^{n}$ of dimension $k$. Any set of linearly-independent rows of a generator matrix $G$ forms a basis of the code ${\cal C}\equiv{\cal C}_{G}$ generated by $G$; the code dimension is $k=\mathop{\rm rank}G$. The dual code ${\cal C}^{\perp}\equiv{\cal C}_{G}^{\perp}$ of dimension nkn-k is formed by all vectors in $F^{n}$ orthogonal to non-zero vectors in $\cal C$} or, equivalently, to the rows of $G$, a check matrix of the code ${\cal C}_{G}^{\perp}$.
@@
The usual inner product in $F^{\ell}$ is related to the group trace with the help of a linear map[37] $\boldsymbol{\widehat{\ }}:F[G]\to F[G]$,
@@
[…]
@@
$\partial^{\prime\prime\prime}(a\otimes b)\equiv\partial^{\prime}a\otimes b+(-1)^{i}a\otimes\partial^{\prime\prime}b,$ (11)
@@
[…]
@@ III.1 Multi-block complexes and associated quantum codes
With Eq. (20), it is easy to verify that the resulting matrices contain only three blocks, $A\equiv A_{1}\otimes I_{B}\otimes I_{C}$, $B\equiv I_{A}\otimes B_{1}\otimes I_{C}$, and $C\equiv I_{A}\otimes I_{B}\otimes C_{1}$, and can be explicitly written as
@@
$R_{i}$ $=$ $\left[\begin{array}[]{cc}Q_{i-1},&0\\ (-1)^{i-1}I({m_{i-1}})\otimes N,&Q_{i}\end{array}\right],\;\,1<i\leq D,\qquad$ (36)
@@
For a $D$-dimensional MBC complex with boundary matrices $Q_{j}$, $1\leq j\leq D$, it is easy to verify that all $D$ blocks are used in an equivalent fashion, i.e., all chain complexes constructed from the same $D$ commuting blocks taken in different orders are equivalent (can be obtained from each other by block row and column permutations). In particular, matrix $Q_{j}$ has $j$ non-zero blocks in each column and D+1jD+1-j non-zero blocks in each row. Furthermore, in each pair of matrices $Q_{j}$ and $Q_{j+1}^{T}$ with $m_{j}={D\choose j}$ column blocks, $1\leq j<D$, each pair of matching column-blocks contains a full set of blocks $\{A_{i}\}$, exactly $j$ such blocks in $Q_{j}$ and the remaining DjD-j (transposed) blocks in $Q_{j+1}^{T}$.
@@
$\mathop{\mathrm{M}}\nolimits(a)]_{\alpha,\beta}\equiv\sum_{g\in G}a_{g}\delta_{\alpha,g\beta},$ (58)
where group elements $\alpha,\beta\in G$ are used to index rows and columns, and $\delta_{\alpha,\beta}=1$ if $\alpha=\beta$ and 0 otherwise is the Kronecker delta. It is easy to verify that for group elements $g\in G$, matrices $\mathop{\mathrm{M}}\nolimits(g)$ form the regular $F$-linear representation of $G$.
@@ III.3 Equivalent AMC codes
Therefore, we can always ‘‘rescale’’ the elements in the original set, and choose $a_{j}=1+\ldots$, $j\leq D$. Further, we can try to construct a $D$-local complex in $\mathbb{Z}^{D}$, e.g., by choosing group generators in the support of each element, e.g., $a_{j}=1+x_{j}+(\text{other monomials})$. Clearly, the second transformation changes the group representation and the basis vectors associated with the lattice.
For example, with weight-two elements $a_{i}=1+x_{i}$, $1\leq i\leq D$, we get a complex locally equivalent to $D$-dimensional QHP codes. Assuming the entire group can be generated this way [^3] Otherwise, the spaces in an AMC complex can be decomposed onto a direct sum of subspaces corresponding to cosets of the subgroup $G_{X}=\langle\{x_{j}\}\rangle$ in the original group $G$; see Sec. IV-C in Ref. 28. In the abelian case these complexes are permutation-equivalent to a complex in the subgroup $G_{X}\leq G$.,
@@
we get a $D$-complex locally equivalent to $D$-dimensional toric codes. With a quantum CSS code associated with level-$j$ subspace, for errors with irreducible weight smaller than the injectivity radius $R$, this guarantees identical syndrome weights, that is, identical confinement profiles[22] up to $R$. We also expect the code distance to grow as $d={\cal O}(R^{\alpha})$, with the exponent $\alpha=\min(j,D-j)$ expected for level-$j$ $D$-dimensional toric codes, see Statement 4 below. Further, we can use a measuring circuit designed for a regular level-$j$ $D$-dimensional toric code: its validity is guaranteed as a local property. [^4] Note, however, that the circuit distance need not always coincide with that of the original code. As an example, while any single-ancilla measurement circuit would work for a toric code[59], rotated surface codes require a carefully designed $\sf Z$-$\sf N$ addressing scheme for fault-tolerance[51].
@@
As an immediate consequence, the corresponding homology groups have orders $\left|H_{j}({\cal Q})\right|={D\choose j}\kappa$. Previously, the same or equivalent results were derived for the special case of $D=2$ AMC codes (quasi-cyclic GB codes), see Refs. [24, 44]. Also note a similarity with the corresponding equations for $D$-dimensional QHP codes. The main difference with Eqs. (18) and (19) is the linear scaling with the group order $\ell$, which is why the AMC construction gives much smaller codes.
@@
Let $G$ be a finite abelian group presented with the list of generators $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{m})$, and $K\equiv G^{\times D}$ its $D$-fold tensor product with itself, with the corresponding list of generators $\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{D}\}$. Given a checkable[37] abelian group-algebra code in $F[G]$ of length $n$, dimension $k$, and distance $d$, with the orthogonal element $a\equiv a(\boldsymbol{x})$, consider the group algebra elements $a_{i}\equiv a(\boldsymbol{x}_{i})\in F[K]$, $1\leq i\leq D$, and the corresponding $D$-complex ${\cal Q}=\mathop{\text{\small\sc AMC}}\nolimits(a_{1},a_{2},\ldots,a_{D})$. Its $j$-th subspace has dimension $n_{j}\equiv\dim({\cal Q}_{j})={D\choose j}n^{D}$, while the corresponding homology and co-homology groups have the same dimension $k_{j}={D\choose j}k^{D}$ and homological distances $d_{j}=d^{j}$ and $\tilde{d}_{j}=d^{D-j}$, respectively.
In the special case of a symmetric even-$D$ AMC code at level $j=D/2$, this gives the encoding rate $k_{j}/n_{j}=(k/n)^{D}$ and the distance $d^{D/2}$. Thus, a finite rate and a linear distance scaling in the original code family would give a finite rate and the distance scaling as a square root of the code length, similar to GB codes equivalent to QHP codes. While we are not aware of a proof that good linear abelian group-algebra codes exist (known results include good code families based on non-abelian dihedral groups $D_{2m}$, random quasi-abelian codes [46], and non-linear cyclic codes[47] that attain the Gilbert-Varshamov bound), already the classical results on the cyclic codes[48, 49] are sufficient to show the existence of AMC codes with finite relative distances and power-law distance scaling, with the exponent arbitrarily close to $1/2$.
@@
Another parameter relevant in a fault-tolerant setting is the syndrome distance $d_{\rm S}$, the minimum weight of a non-zero syndrome vector. An obvious upper bound is the minimum column weight in the check matrix. A boundary operator $Q_{j}$ in an AMC $D$-complex $\cal Q$} has $j$ or D+1jD+1-j non-zero blocks in each block column or row, respectively. If we use group algebra elements of the same weight $w$, this gives uniform column weights $jw$ and row weights (D+1j)w(D+1-j)w. For a code $\mathop{\rm CSS}\nolimits(Q_{j},Q_{j+1}^{T})$ in the space ${\cal Q}_{j}$ of such a $D$-complex, this gives a simple upper bound on the syndrome distance,
@@ IV.1 Family of $4$-dimensional rotated toric codes
Specifically, we constructed such codes for $7\leq\ell\leq 30$ and corresponding block lengths $n=6\ell$ (four distinct polynomials cannot be found for $2\leq\ell<7$). Generating polynomials and parameters of smallest codes for each minimum distance $d$ are listed in Tab. 1. All constructed codes have the same dimension $k=6$ and the same characteristic polynomial $h(x)=1+x$. For comparison, in Tab. 1 we also list parameters of two conventional $4D$ toric codes, which require larger block lengths to achieve the same distances.
In addition to conventional code parameters and the defining polynomials, Tab. 1 shows the confinement profile[22], a sequence of minimum syndrome weights corresponding to errors of increasing irreducible weights $w\geq 1$. These are important for single-shot error correction in the presence of syndrome measurement errors. Indeed, for a conventional level-2 4D toric code, the undetectable errors are closed surfaces formed by plaquettes, while such a surface with a boundary is detectable, where edges along the surface boundary correspond to flipped syndrome bits. With the lattice of size $L^{4}$, the perimeter is expected to scale as $\mathcal{O}(w^{1/2})$, for error weights $w\leq w_{\rm max}=\mathcal{O}(L^{2})$, with the first five terms in the large-$L$ sequence $4,6,6,8,8$. The confinement profiles of the codes in Tab. 1 (computed using the dist_m4ri program [50]) follow this sequence, as long as the corresponding minimum-perimeter configuration does not contain a pair of edges identified by a lattice vector. This is substantially different from $3+3$ GB codes previously studied in Ref. 25, where for all designed codes the confinement profiles (starting with $3,4,3$) are increasing only for irreducible error weights $w\leq 2$.
@@
Table 1: Smallest $D=4$ AMC2 qubit codes (rotated $4D$ toric codes) with increasing distances $d$, constructed from cyclic groups $C_{\ell}$ with $\ell\leq 30$ (above the line) and two conventional $4D$ toric codes (below the line). Confinement profile [22] (‘‘confin’’) column lists the minimum syndrome weights for irreducible errors of weights from $1$ to $5$, except for codes with $\ell=7$ and $\ell=2^{4}$ where the biggest irreducible errors have weights $4$ and $3$, respectively.
@@ IV.2 Circuit design
As an example, after removing the first and the last row blocks, respectively, from the CSS matrices $H_{X}=R_{2}$ and $H_{Z}=R_{3}^{T}$ in Eqs. (45) and (52), we get in the case of a binary field,
@@
where the subscripts next to block names label the columns and the circumflex ‘‘hat’’ accent indicates transposed blocks. Focusing on bigger $3\times 3$ blocks, we have a two-block structure similar to that in lifted-product or GB/2BGA codes, with two and four ("2+42+4") non-zero entries in each subrow. The circuits designed[25] for $2+4$ 2BGA/GB codes, a generalization of the И-Z addressing scheme[51] for rotated surface codes, also work in the present case. Namely, a valid time-optimal circuit is obtained if we address the qubits corresponding to the two terms in the $D$ blocks at steps 1 and 6 (first and last), with the opposite order for $H_{X}$ and $H_{Z}$, while addressing the qubits from the remaining blocks in the intervening time steps. Specifically, we used the following addressing scheme:
@@
Here each column corresponds to a circuit time step, and each row to a block row of the matrices (67), (71), with secondary superscripts labeling the location corresponding to a free term (‘‘0’’) and the other monomial (‘‘1’’). Circuits corresponding to the original matrices with 2nd, 3rd, or 4th row removed were constructed identically, with the help of column permutations rendering the matrices in the form (67), (71), up to block labels.
[…]
@@
For accurate decoding, one needs to know probabilities of individual faults in the circuit and their effects on the detectors and the logical observables. This information for a given circuit is computed automatically by Stim, which also aggregates the faults producing the identical effect, and can be exported as a detector error model (DEM) file[53]. Information about aggregated faults in a DEM file can be regarded as columns of stabilizer-generator $H_{X}$ and logical $L_{X}$ matrices (with known aggregated fault probability for each column) for an asymmetric quantum CSS code. In particular, the distance $d_{Z}$ of this code is identified as the circuit distance. We used vecdec program [54] to export the matrices from the DEM files associated with individual circuits, and dist_m4ri program to compute the corresponding distances (with the distances $d\leq 6$ computed exactly with a greedy enumeration algorithm, and upper bounds for larger distances using a random information set algorithm[55], exact with a high probability).
@@
Circuit simulations have been done with Pauli-frame circuit simulator Stim [53] written by Craig Gidney, and the decoding using vecdec program [54]. Specifically, before attempting belief-propagation (BP) decoding, the program tries to match the syndrome data against a pre-computed list of small-weight error clusters; such a simple pre-decoder gives a substantial acceleration, especially at smaller error probabilities. If pre-decoding fails, the corresponding clusters are discarded, and a BP decoder with serial schedule, using both average and instantaneous logarithmic likelihood ratios (whichever converges first) is engaged, with up to 50 steps. Finally, if that fails, the logarithmic likelihood ratios (LLRs) are transmitted to the OSD-1 decoder. A sample of circuit simulation data for codes $[[42,6,4]]$, $[66,6,6]]$, and $[[96,6,8]]$ (see Tab. 1), with the standard error model described earlier and the ‘‘1212’’ syndrome-addressing cycle, is shown on a threshold-style plot in Fig. 1. To speed up the decoding, we simplified the DEM by dropping all ‘‘minority’’ detector events, which somewhat increases logical error rates [^6] This is unlike with the BP+OSD package[60, 61] whose performance may actually be degraded by additional detector events, see in Appendix A of Ref. 62. In comparison, for the decoder used here, removing minority detector events increases logical error rates, e.g., from $0.20\%$ to $0.36\%$ at $p=0.1\%$ with the $[[42,6,4]]$ code and the “1212” circuits.. Nevertheless, asymptotic slopes at small $p$ consistent with $d/2$ (where $d$ is the code distance) is in agreement with the circuit distances computed independently. The (pseudo)threshold position (crossing point) is at $p_{c}\approx 1.1\%$, in approximate agreement with the circuit simulation results for similar codes in Ref. 34. We note that our value is (marginally) higher than the values for toric and surface codes under the same error models.
@@
Figure 1: Logical error rate as a function of the circuit noise parameter $p$, using the syndrome-addressing scheme ‘‘1212’’ and $Z$-basis final measurements. For faster simulations, only $Z$-measurement detector events have been included. Results for short even-distance codes $[[42,6,4]]$, $[66,6,6]]$, and $[[96,6,8]]$ (see Tab. 1) are shown. Inset shows the success probability $1-p_{L}$ as a function of $p$ in the vicinity of the crossing point(s) at $p\approx 1.1\%$
@@
In addition, to verify the ‘‘single-shot’’ properties of the constructed codes, we have tried sequential one-step sliding window decoding[18, 57, 23], with windows of different sizes. Details of our implementation are given in Ref. 25. Some simulation results for the same three codes as in Fig. 1 are shown in Fig. 2, where the logical error rates $p_{L}$ are plotted as a function of the decoding window size $T$ at a fixed circuit error rate, $p=10^{-3}$. Up to $5\times 10^{6}$ samples have been generated for each point, resulting in very small statistical errors.
For the data in Fig. 2 we used full DEMs for decoding, including the detector events associated with both $X$- and $Z$-measurements, which resulted in somewhat lower $p_{L}$ at full-block decoding ($T=9$) than those in Fig. 1 at the same $p$. For each code, we tried all three syndrome measurement cycles. Simplest is the ‘‘1111’’ cycle, where the same three row blocks are measured in each round. While discarding most of the redundant generators, this cycle is the most accurate with full-block decoding. In comparison, the ‘‘1234’’ cycle which addresses most uniformly all low-weight stabilizer generators, gives the lowest fail rates with single-shot ($T=1$) and two-shot ($T=2$) decoding protocols. Note that only during the rounds labeled with ‘‘1’’ a full-rank set of stabilizer generators is measured.
Figure 2: Logical error rates computed with full detector error models (both $X$ and $Z$ detector events) for the same codes as in Fig. 1, plotted as a function of the decoding window size $T$. Shading indicates the statistical error at one standard deviation. Here $T=9$ corresponds to (most accurate) full-block decoding, while $T<9$ correspond to various one-step sliding window decoding protocols. The data labeled as ‘‘1111’’, ‘‘1212’’, and ‘‘1234’’ correspond to different syndrome measurement cycles as explained in the text. At larger $T$, syndrome measurement errors are less important, and the ‘‘1111’’ cycle gives lower error rates due to shorter measurement cycle. In contrast, with single- ($T=1$) and two-shot ($T=2$) decoding, the error rates are lower with the other two cycles which better utilize the syndrome redundancy. In comparison, in the ‘‘1111’’ measurement cycle, most of the redundant checks are discarded.